OpenStudy (anonymous):

Evaluate0.5g-g2 when g=4 -14 4 18 -6

OpenStudy (anonymous):

Wherever there is a g, put in a 4 so 0.5g - g2 becomes 0.5*4 - 42 0.5*4 - 42 = 2 - 16 = -14